OpenStudy (anonymous):

help me solve (x+3)(2x-1)

OpenStudy (anonymous):

x+3=0 x=-3 and 2x-1=0 x=1/2

OpenStudy (anonymous):

im sorry i forgot to say multiply

OpenStudy (anonymous):

2x^2+5x-3

OpenStudy (anonymous):

thank you